# A Comprehensive Survey on Kolmogorov Arnold Networks (KAN)

### v4

@@ 6.2 Time Series Prediction
Kolmogorov-Arnold Networks (KAN) have shown significant potential in time series prediction. Time series prediction involves using past data to forecast future values and has wide applications in fields such as financial market analysis, weather forecasting, and medical monitoring. KAN, with its unique architecture and flexible activation functions, can capture complex patterns and both short-term and long-term dependencies in time series data.
6.2.1 T-KAN and MT-KAN
T-KAN (Time-Kolmogorov-Arnold Network)
The T-KAN model is specifically designed for handling time series data. It leverages the flexibility of KAN by using learnable univariate activation functions to capture nonlinear relationships and complex patterns in time series. Compared to the basic KAN model, T-KAN achieves higher performance through the following improvements:
  1. 1.
    Dynamic Activation Functions: Traditional neural networks typically use fixed activation functions, such as ReLU or Sigmoid, while T-KAN uses learnable univariate activation functions. This allows the model to dynamically adjust its activation patterns, better adapting to variations in time series data [5].
  2. 2.
    Capturing Nonlinear Relationships: T-KAN is more effective at capturing nonlinear relationships in time series data, excelling in tasks such as financial market prediction and energy consumption forecasting. For instance, Xu et al. (2024) demonstrated T-KAN’s superior performance in stock market prediction by effectively modeling historical price and trading volume data, achieving accurate short-term price forecasts.
The innovation of T-KAN lies in its ability to better adapt to nonlinear changes in time series data through learnable activation functions. This flexibility makes T-KAN particularly effective at capturing complex temporal dependencies. Additionally, compared to the basic KAN model, T-KAN optimizes the learning mechanism of activation functions, enabling more precise reflection of dynamic changes in time series data.
MT-KAN (Multi-Task Kolmogorov-Arnold Network)
The MT-KAN model extends the capabilities of T-KAN to multi-task learning, enabling it to handle multiple related tasks simultaneously. Compared to the basic KAN model and T-KAN, MT-KAN achieves improvements through the following methods:
  1. 1.
    Shared Network Structure: MT-KAN enhances overall prediction performance by sharing network layers and parameters, leveraging inter-task relationships. For example, in multi-task time series forecasting tasks such as power load forecasting and air quality monitoring, MT-KAN achieves higher prediction accuracy by modeling the interactions in multivariate time series.
  2. 2.
    Parameter Sharing and Task Interaction: By sharing certain model parameters, MT-KAN can transfer useful feature representations between tasks, improving the prediction performance of each task. This approach not only enhances model accuracy but also reduces the amount of training data required.
6.2.2 SigKAN
SigKAN (Signature-Weighted Kolmogorov-Arnold Network)
SigKAN enhances KAN’s performance in time series prediction by incorporating path signatures. Path signatures capture the essential geometric features of time series paths. By integrating this information with KAN’s output, SigKAN can provide a more comprehensive representation of temporal data. Specifically:
  1. 1.
    Path Signature Integration: Path signatures are tools that capture the geometric features of time series data. By integrating path signatures, SigKAN can better understand and predict complex time series patterns.
  2. 2.
    Dynamic Adaptation: SigKAN dynamically adapts to different data patterns by learning adjustable parameters of path signatures, improving its performance in multi-task time series forecasting. Inzirillo and Genet (2024) suggested that SigKAN outperforms traditional methods in multi-task time series forecasting [17].
6.2.3 Experiments and Results
In several experiments, T-KAN, MT-KAN, and SigKAN demonstrated their excellent performance in time series prediction. For example, experiments on financial time series data showed that T-KAN and MT-KAN excel in predicting stock implied volatility, significantly outperforming traditional MLP and RNN models. These results indicate that KAN models can achieve high-precision forecasts even in complex and highly volatile financial data environments. SigKAN, by combining the advantages of path signatures and KAN, further enhances performance in multi-task time series forecasting, showing superior performance in tasks such as market trading volume prediction and absolute return prediction.
Research by Kundu et al. (2024) also demonstrated the superior performance of KAN in satellite traffic prediction. Experiments showed that KAN models outperform traditional MLP models when dealing with complex satellite traffic data. KAN not only surpasses MLP in prediction accuracy but also demonstrates excellent parameter efficiency, achieving higher prediction performance with fewer trainable parameters [18].
KAN has shown significant advantages in time series prediction, effectively capturing complex temporal dependencies and nonlinear patterns through its unique architecture and flexible activation functions. These characteristics give KAN broad application potential and excellent performance in fields such as financial forecasting and energy management.
6.3 Graph-Structured Data Processing
Kolmogorov-Arnold Networks (KAN) have demonstrated significant advantages in processing graph-structured data, particularly in the fields of graph learning and Graph Neural Networks (GNNs). Graph-structured data is prevalent in various domains such as social networks, molecular structures, and biological networks. Processing these data requires powerful models to capture complex relationships and structural information between nodes. Graph-structured data processing faces several challenges, including:
  1. 1.
    Complex Node Relationships: The relationships and dependencies between nodes and edges in graph-structured data are complex, making it difficult for traditional neural networks to handle these structures directly.
  2. 2.
    High-Dimensional and Large-Scale Graphs: Graph-structured data often have high-dimensional features and large-scale structures, imposing high demands on computational resources and model scalability.
  3. 3.
    Heterogeneity and Diversity: The structure and features of different graphs can vary significantly, requiring models to have strong generalization capabilities to adapt to various graph data.
6.3.1 Graph Kolmogorov-Arnold Networks (GKAN)
GKAN extends the principles of KAN to graph-structured data. By introducing learnable univariate functions on the edges of the graph, GKAN innovatively changes the way information is processed in graph structures, replacing the fixed convolution architecture used in traditional GCNs.
GKAN: Graph Kolmogorov-Arnold Networks
GKAN’s architecture demonstrates innovation and flexibility when processing graph-structured data. Its main features include:
  1. 1.
    Node Feature Aggregation: GKAN employs two main architectures differing in the sequence of aggregation and activation function application. The first architecture aggregates node features (e.g., summation or averaging) before processing through the KAN layer. This method emphasizes applying learnable activation functions after aggregation, allowing the model to better capture complex relationships within local graph structures [20].
  2. 2.
    Embedding Layer Processing: The second architecture applies KAN layers between node embeddings at each layer before aggregation. By introducing learnable univariate activation functions between layers, GKAN can flexibly adapt to feature changes across different graph structures. This approach allows the model to dynamically adjust its representational capacity between different layers, enhancing its effectiveness in processing complex graph data [20].
  3. 3.
    Dynamic Adaptability: GKAN can dynamically adjust according to the structure and features of different graphs, enhancing its adaptability and robustness. This characteristic makes GKAN particularly effective in handling heterogeneous and diverse graph data. Compared to the basic KAN model, GKAN exhibits greater scalability and adaptability when processing large-scale and high-dimensional graph-structured data [5, 18].
Experiments and Results
In multiple experiments, GKAN demonstrated excellent performance in processing graph-structured data. For example, experiments on the Cora dataset showed that GKAN significantly outperforms traditional GCN models in node classification tasks. Specifically, with 100 features, the accuracy of traditional GCN is 53.5
Additionally, GKAN’s performance on other graph datasets also indicates its superiority in handling large-scale and high-dimensional graph-structured data, especially in tasks such as social network analysis and molecular graph analysis, demonstrating its outstanding performance and broad application potential.
6.4 Hyperspectral Image Classification
Kolmogorov-Arnold Networks (KAN) have demonstrated significant advantages in hyperspectral image (HSI) classification, particularly in handling high-dimensional data and complex spatial-spectral correlations. Hyperspectral images contain rich spectral information, which can be used to distinguish different materials, making them widely applicable in fields such as agriculture, environmental monitoring, and military applications. Hyperspectral image classification faces several challenges, including:
  1. 1.
    High-Dimensional Data: Hyperspectral images typically contain hundreds of spectral bands, requiring effective dimensionality reduction and feature extraction methods [21].
  2. 2.
    Complex Spatial-Spectral Correlations: Each pixel in a hyperspectral image contains both spectral and spatial information, increasing the complexity of classification [22].
  3. 3.
    Data Sparsity: In many applications, annotated hyperspectral image data are very limited, making effective classification with small sample sizes a significant challenge [21].
6.4.1 Advantages of KAN in Hyperspectral Image Classification
Kolmogorov-Arnold Networks (KAN) can effectively address these challenges in hyperspectral image classification through their unique architecture and flexible activation functions. Specifically:
Wavelet-based KAN (Wav-KAN)
Wav-KAN introduces wavelet functions as learnable activation functions into the KAN architecture to achieve nonlinear mapping of input spectral features. This method can effectively capture multi-scale spatial-spectral patterns through dilation and translation, significantly enhancing classification performance. For example, Seydi et al. (2024) demonstrated Wav-KAN’s superior performance on benchmark datasets such as Salinas, Pavia, and Indian Pines, with significant improvements in classification accuracy compared to traditional MLP and Spline-KAN models [21].
SpectralKAN
SpectralKAN uses B-spline functions as activation functions, combining these functions to represent multivariate continuous functions for feature extraction and classification of hyperspectral images. SpectralKAN introduces a spatial-spectral KAN encoder, where the spatial KAN encoder extracts spatial features and compresses the spatial dimensions, while the spectral KAN encoder extracts spectral features and classifies them into changing and non-changing categories. This method not only improves classification accuracy but also reduces model parameters, FLOPs, GPU memory, and training and testing time, enhancing computational efficiency [22].
Detailed Analysis
  1. 1.
    High-Dimensional Data Processing: Wav-KAN and SpectralKAN effectively handle the high-dimensional data problem in hyperspectral images by using learnable activation functions. Wav-KAN uses wavelet functions to perform multi-scale decomposition of spectral data, capturing features at different scales, enabling the classification model to comprehensively understand the spatial and spectral information in the data [21].
  2. 2.
    Spatial-Spectral Correlations: SpectralKAN captures complex spatial-spectral correlations in hyperspectral images by introducing a spatial-spectral KAN encoder that separately processes spatial and spectral features. The spatial KAN encoder first extracts spatial features and performs dimensionality reduction, and then the spectral KAN encoder further extracts spectral features and classifies them. This two-stage processing method significantly improves the model’s ability to capture complex spatial-spectral patterns [22].
  3. 3.
    Data Sparsity Issue: By reducing model parameters and improving computational efficiency, SpectralKAN demonstrates strong adaptability in handling small sample hyperspectral image data. Experimental results show that SpectralKAN outperforms traditional methods on multiple datasets, maintaining high classification accuracy even with limited computational resources [22].
Experiments and Results
In multiple experiments, Wav-KAN and SpectralKAN demonstrated excellent performance in hyperspectral image classification. For example, Wav-KAN’s experimental results on the Salinas dataset showed that this method can better capture complex spatial-spectral correlations, significantly improving classification accuracy [21]. SpectralKAN’s experimental results on multiple HSI-CD datasets also showed that this method significantly reduces computational resource requirements while maintaining high classification accuracy [22].
6.5 Quantum Architecture Search
Kolmogorov-Arnold Networks (KAN) have demonstrated significant potential in Quantum Architecture Search (QAS), which aims to optimize circuit design in quantum computing to enhance the efficiency and performance of quantum algorithms. By leveraging KAN’s flexible architecture and interpretability, researchers can capture and optimize key patterns and relationships in quantum computing.
6.5.1 VQ-KAN (Quantum Variational KAN)
The VQ-KAN model is specifically designed for quantum computing optimization problems, utilizing KAN’s flexibility by employing learnable univariate activation functions to capture and optimize the complex relationships between quantum states and operations. Compared to traditional quantum algorithm optimization methods, VQ-KAN improves performance through the following innovations:
  1. 1.
    Dynamic Activation Functions: Traditional quantum optimization algorithms rely on fixed parameters and strategies, whereas VQ-KAN uses learnable univariate activation functions that can dynamically adjust their activation patterns to better adapt to different quantum operations and states [23].
  2. 2.
    Capturing Nonlinear Relationships: VQ-KAN is more effective at capturing nonlinear relationships in quantum operations, excelling in complex quantum circuit optimization tasks. Research shows that VQ-KAN has significant advantages in optimizing the performance and resource utilization of quantum algorithms [23].
Quantum State Representation
By introducing path signatures, VQ-KAN further enhances its performance in quantum state representation. Path signatures capture the geometric features of quantum states, and by combining this information with KAN’s output, VQ-KAN can represent quantum states more comprehensively. This method not only improves representation accuracy but also enhances the model’s robustness and adaptability when handling different quantum states and operations [23].
Experiments and Results
In multiple experiments, VQ-KAN demonstrated excellent performance in quantum architecture search. For example, experiments in quantum circuit optimization showed that VQ-KAN could better capture and optimize complex quantum operations, improving the performance and resource utilization of quantum algorithms [23].
6.6 Medical Image Segmentation and Generation
Kolmogorov-Arnold Networks (KAN) have shown significant advantages in medical image segmentation and generation, particularly in improving segmentation accuracy and generating high-quality images. KAN, through its unique architecture and flexible activation functions, can capture complex spatial-spectral correlations, optimizing medical image processing tasks.
6.6.1 Application of U-KAN in Medical Image Segmentation
U-KAN integrates KAN layers into the U-Net architecture, enhancing the accuracy and interpretability of segmentation models. Specifically:
  1. 1.
    U-KAN (U-Net with Kolmogorov-Arnold Network): U-KAN integrates KAN layers into the U-Net architecture, using learnable univariate activation functions to replace traditional fixed activation functions. This method captures complex nonlinear patterns in input features, improving segmentation model performance. For example, Li et al. (2024) demonstrated U-KAN’s superior performance in medical image segmentation, achieving high accuracy in various medical image segmentation benchmarks [24].
  2. 2.
    U-KAN in Generation Tasks: U-KAN is also applied in noise prediction for diffusion models, improving high-quality image generation by enhancing the U-Net noise predictor. This method shows U-KAN’s potential in generation tasks, capable of producing high-quality task-oriented model architectures [24].
Detailed Analysis
  1. 1.
    Capturing Complex Nonlinear Patterns: U-KAN, through its flexible activation functions, can effectively capture and represent complex nonlinear patterns in medical images, which is crucial for improving image segmentation accuracy and reliability [24].
  2. 2.
    Model Interpretability: Since KAN’s activation functions are learnable, U-KAN models are more interpretable. This allows medical professionals to better understand and trust segmentation results, especially in critical medical diagnoses [24].
Experiments and Results
In multiple experiments, U-KAN demonstrated excellent performance in medical image segmentation and generation. For example, experimental results on various medical image segmentation benchmarks showed that U-KAN significantly improved segmentation accuracy and remained efficient even with limited computational resources [24].
6.7 Word-Level Explainable Language Model
Kolmogorov-Arnold Networks (KAN) have shown potential in natural language processing, particularly in word-level explainable language models. These models aim to provide more transparent and interpretable language models by explaining the meaning and relationships of each word in a sentence.
6.7.1 UKAN (Understandable KAN)
UKAN leverages KAN’s interpretability to construct an explainable language model, where specific neurons encode individual words, and neuron activations can be fully interpreted as the basis of words. To this end, UKAN proposes a continuous word2vec model, where the meaning of a word is represented by a continuous spectrum of its distance from a word base.
  1. 1.
    Continuous Word2vec Model: By representing word meaning as a continuous spectrum of its distance from words in the word base, UKAN provides finer-grained and interpretable word representations in natural language processing [23].
  2. 2.
    Word-Level Explanation: In the UKAN model, the activation of each word can be fully explained by its position in the word base. This method not only enhances model interpretability but also makes the language model’s prediction and decision-making process more transparent and intuitive [23].
Symbolic Regression and Logical Programming
UKAN combines symbolic regression and logical programming techniques to achieve higher interpretability and accuracy in processing complex semantic relationships and reasoning tasks. For example, by utilizing symbolic regression techniques, UKAN can dynamically adjust activation functions to better capture nonlinear relationships between words, enhancing the model’s performance in natural language understanding tasks [23].
Experiments and Results
In multiple experiments, UKAN demonstrated excellent performance in natural language processing tasks. For example, experimental results on language model tasks showed that UKAN’s word-level explainable language model provided higher interpretability and transparency, significantly improving prediction accuracy and user trust [23].

### v5

@@ 6.1 Time Series
Kolmogorov-Arnold Networks (KAN) have shown significant potential in time series prediction. Time series prediction entails utilizing historical data to forecast future values and has extensive applications in areas such as financial market analysis, weather forecasting, and medical monitoring. KAN, with its unique architecture and flexible activation functions, can capture complex patterns and both short-term and long-term dependencies in time series data.
TKAN (Temporal Kolmogorov-Arnold Networks)
T-KAN constructed by Genet and Inzirillo [25] is combined of RKAN (Recurrent Kolmogorov-Arnold) Network layers and Gating Mecanism. The special structure enables T-KAN to solve the complex time series problems and maintain long memory. In general, it combined RKAN and slightly modified LSTM, and its activation function can help capture complex nonlinearities.
T-KAN and MT-KAN (Time Series Kolmogorov-Arnold Networks and Multivariate Time Series Kolmogorov-Arnold Networks)
Xu et al (2024) [26] propose T-KAN and MT-KAN model to solve time series prediction problems. T-KAN employs a two-layer (2-depth) network structure where each layer consists of spline-parametrized univariate functions. It is designed to handle univariate time series data. MT-KAN is a variant of T-KAN, which is designed to handle multivariate time series data. MT-KAN employs mechanisms to model cross-variable interactions, which allows it to exploit the relationships among different variables.
SigKAN (Signature-Weighted Kolmogorov-Arnold Networks)
SigKAN is composed of SigLayer and KAN layer. Specially, the KAN layer is modified by Gated Residual Networks (GRN), which makes the model modulate the information flow during the learning task. The Learnable Path Signature layer (SigLayer) makes it possible to compute the path signatures for each path with learnable scaling. SigKAN can better understand and predict complex time series patterns and shows great dynamic adaptation [27]
WormKAN
Xu et al (2024) [28] applied KAN in decoder and encoder, with a self-representation layer implemented as a 2-layer KAN. Thy compared the model with StreamScope, TICC, AutoPlait. The results show that WormKAN has higher F1 score and ARI. It shows potential in co-evolving time series tasks.
HiPPO-KAN
Lee et al [29] integrate High-order Polynomial Projection (HiPPO) theory into the Kolmogorov-Arnold network (KAN) framework. They highlight that the model can help address the problem of time series forecasting in the context of cryptocurrency markets, specifically focusing on the BTC-USDT trading pair. Compared with other models, HiPPO-KAN with 500 window size has a smaller MAE, HiPPO-KAN with 1200 window size has a smaller MSE, which shows a better accuracy in time series tasks.
KAN-AD
. Kolmogorov–Arnold Net-works (KAN) has potential to decompose the complex temporal sequences into different functions, which makes it easy to be controlled when training. However, this method may be influenced by local anomalies. Zhou et al [30] use the Fourier series to emphasize global temporal patterns, thereby mitigating the influence oflocal peaks and drops. Their evaluation shows that KAN-AD has a better efficiency and costs less time compared with state-of-art technologies.
6.2 Interpretability and Symbolic Regression
Kolmogorov-Arnold Networks (KAN) provide enhanced interpretability compared to traditional neural networks, such as Multi-Layer Perceptrons (MLP)[31]. In this section, we will introduce some explainable models.
Monotonic Kolmogorov-Arnold Network (MonoKAN)
Polo-Molina et al [32] construe a new structure of KAN which can enhance the model’s interpretability. They replace the B-splines with cubic Hermite splines which has well-known sufficient conditions for monotonicity. experiments and mathematical reasoning demonstrated monotonicity. Since the momotonicity of the model is proved, the new structure is thought to be more explainable.
KAN2CD
KAN2CD [33] is designed to enhance interpretability in two manners. The first method is to replace the MLP layer in CMD (cognitive diagnosis models). The second method uses KAN to deal with student embedding, exercise embedding and concept embedding and then the output will be combined and learned by KAN. Yang et al highlight that the learned structures of KAN make KAN2CD more explainable than traditional CMDs.
Bayesian Kolmogorov Arnold Networks (BKANs)
Hassan [34] combines KANs with Bayesian inference and then tests the new model on 2 medical datasets. The results show that BKANs show great improvement in interpretability and accuracy compared with traditional deep learning models.
CoxKAN
Knottenbelt et al [35] combine Cox Proportional Hazards Model and Kolmogorov Arnold Network for interpretable, high performance survival analysis, named as CoxKAN. Then they evaluate CoxKAN on 4 synthetic datasets and 9 real medical datasets. The results on 9 real medical datasets show that CoxKAN always has a better performance than original Cox Proportional Hazards Model. Moreover, CoxKAN can obtain symbolic formulas for the hazard function and visualizing KAN activation functions and visualize the activation functions, which show great interpretability.
KAN-XGBoost
Amouri et al [36] introduce hybrid model consists of KAN and XGBoost algorithm. The hybrid model integrates the learnable function of KAN and high accuracy in classification task, show great accuracy and interpretability in the Intrusion Detection System. It has higher higher accuracy, precision, recall and F1 score on N-BaIoT dataset compared with standalone MLP and KAN networks.
7 Health and Medical Field
U-KAN
Li et al [li2024ukanmakesstrongbackbone] integrate KAN into the tokenization process after using Convolution Phrase to abstract features to enhance U-Net. Then, they extend the model to diffusion models. Their test on medical datasets BUSI and GlaS shows that U-KAN has a better compared with state-of-the-art segmentation models on three heterogeneous medical scenarios. Tang et al [37] apply U-KAN to 3D brain tumor segmentation using multi-modal MRI data including a variant called UKAN-SE.They compare the models with current models like U-Net, attention U-Net and so forth. The result show that U-KAN and UKAN-SE have better efficiency and accuracy with less training time.
KACQ-DCNN (Uncertainty-Aware Interpretable Kolmogorov-Arnold Classical-Quantum Dual-Channel Neural Network )
[…]
8 Computer Vision and Image Processing
KAN-CUT
Mahara et al [39] replace the two-layer MLP with a two-layer KAN in the Contrastive Unparied Image-to-Image Translation (CUT) model. They use some well-known GAN models as baselines, including CycleGAN, MUNIT and so on. Their results show that KAN-CUT has a smaller FID score in Horse2Dog and Cat2Dog datasets. Moreover, they highlight that KAN-CUT can generate higher quality zebra images based on stripes, structure, and color.
KAN-Block
In order to address the problems caused by uneven illumination and noise effects, Ning et al [40] design a KAN-block to introduce KANs into the U-Net structure in the low-light enhancement task. In their experiments, they choose LOLv1, LOLv2, and LSRW datasets and Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity (SSIM) as two full-reference distortion metrics. Additionally, they employ Learned Perceptual Image Block Similarity (LPIPS) and Fréchet Inception Distance (FID) as two perceptual metrics to evaluate the visual quality of the enhanced outcomes. Their results show that their method is able to recover image details and enhance quality. Moreover, their method shows strong generalization and robustness.
Residual Kolmogorov-Arnold Network (RKAN)
Yu, Wu and Gui [41] integrates the Kolmogorov-Arnold Network (KAN) into the CNN framework as a residual element and then postpone RKAN block. Then, they test the model on some famous datasets including CIFAR-10, Food-101. Their results show that RAKN has potential to solve some challenging problems for traditional CNNs, or even Vision Transformers (ViT). However, they also mention the high cost of the model.
9 Graph Neural Network and Collaborative Filtering
FourierKAN-GCF
Xu et al [42] replace the multilayer perceptron (MLP) with Fourier Kolmogorov-Arnold Network (KAN) as a part of the feature transformation during message passing in GCN. They test the model on MOOC and Anazon Video Games datasets and select BPR-MF and BUIR as baseline. The results shown that FourierKAN-GCF outperforms all baselines in two datasets across various metrics. They especially highlight that their method is easier to train and has representation power than MLPs
Graph Kolmogorov–ArnoldNetwork (GKAN)
Carlo[43] extends Kolmogorov-Arnold Networks (KAN) to graph-structured data by introducing learnable functions to replace traditional GCN’s fixed convolution. This flexible approach allows GKAN to adapt dynamically, enhancing its ability to process complex graph data. Experiments show GKAN outperforms traditional GCNs in tasks like node classification, social network, and molecular graph analysis, especially with large-scale and high-dimensional data.
CF-KAN
Park et al [44] introduce KAN to Collaborative filtering (CF) part in recommender systems. They claim that the through the ability of KAN to learn nonlinear functions at the edge level, exhibit enhanced resistance to catastrophic forgetting compared to MLPs. In their experiments, they compare CF-KAN with CF-MLP on some real-word datasets. The results show that KANs outperform CF-MLP, are highly interpretable, scalable, and will be explored further in diverse recommendation domains.
10 Physics-Informed Neural Networks and Engineering Applications
DeepOKAN
Diab et al [45] introduce a new neural operators called DeepOKAN, using KANs rather than the conventional neural network architectures. In general, DeepOKAN has potential in solving various mechanical problems. Then they compare DeepOAKN with DeepONet. They claim that although DeepONet is effective on classical data-driven, DeepOKAN outperforms it.
KINN (Kolmogorov-Arnold-Informed Neural Network)
In order to solve the problem caused by AI for PDEs, Wang et al [46] propose different PDE forms based on KAN instead of MLP, termed Kolmogorov-Arnold-Informed Neural Network (KINN) for solving forward and inverse problems. In their test, they compared KAN and MLP in strong, energy, and inverse forms of PDEs. The results show that KAN has higher accuracy and convergence speed than MLP in most PDE problems. They highlight that KAN is a promising method to solve PDEs. However, they claim that it is important to determine the grid size according to the problem’s complexity and its long training time is still a problem.
PIKAN (Physics-Informed Kolmogorov-Arnold Networks)
Shuai and Li [47] introduce a KAN-based physics-informed neural network (PINN) which is can efficiently and accurately learn dynamics within power systems. They replace the traditional MLP layers with KAN layers. PIKANs achieve a better accuracy in predicting power system dynamics with a smaller network size. The results show that PIKANs in accurately capturing the dynamics of power systems, which demonstrates a the bright feature of KAN in PINN.